In how many ways can someone pick 5 cubes out of 5 red and 5 orange cubes, if at least one cube of each color has to be picked? Thank you.
Answer by sudhanshu_kmr(1152)


Here we have to pick 5 cubes out of 5 red and 5 orange cubes.
No. of ways to pick 5 cubes from total 10 cubes, if no restriction = 10C5

now, there are 2 conditions

a. when all 5 cubes are red, no of ways = 5C5 = 1
b. when all 5 cubes are orange, no of ways = 5C5 = 1
according to question we can not consider these conditions

no. of ways = 10C5 - 2 = 252 - 2 = 250

Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try the problem yourself.